OpenStudy (anonymous):

csc(theta/2)=sin(theta/2) find all solutions. and solutions in [0,2pi]

OpenStudy (freckles):

Let's think about sin(u)=csc(u)

OpenStudy (freckles):

when is sin(u) the same as it's reciprocal csc(u)?

OpenStudy (freckles):

We also want to solve on the interval theta [0,2pi] that is \[0 \le \theta \le 2 \pi \\ \text{dividing all sides by 2 gives } \\ 0 \le \frac{ \theta}{2} \le \pi\] so solve sin(u)=csc(u) on the interval u [0,pi]

OpenStudy (freckles):

once you have found u we will go back in find theta through the relation theta/2=u